Circular No. 6366

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) M. J. Mumma, M. A. DiSanti, N. Dello Russo, and D. X. Xie, Goddard

Space Flight Center, NASA; M. Fomenkova, University of California at San Diego; and K. Magee-Sauer, Rowan College of New Jersey, report the detection of various parent molecules in C/1996 B2 on Mar. 23 and 24 UT, using the CSHELL cryogenic infrared spectrometer at the NASA Infrared Telescope Facility on Mauna Kea: “A partial listing of the species, the vibrational bands, and the detected lines follows: H_{2}O [111-100, 8 ortholines, and 5 para-lines], HDO [ν_{1}, 3 lines], CO [1-0, 8 lines, J = 0-6], CH_{4} [ν_{3}, 2 lines, R0 and R^{−}(1)], CH_{3}OH [ν_{3}, Q-branch, resolved], C_{2}H_{6} [ν_{7}, rQ_{0}- and ^{r}Q_{1}-branches], HCN [ν_{3}, 7 lines, J = 1-6], and OCS [ν_{1} band, 5 lines]. NH_{3} is provisionally detected in the a- and s-Q-branches (ν_{1}). C_{2}H_{2}, H_{2}CO, and HC_{3}N were searched for but not detected. Numerous unidentified lines are present. All lines appear peaked on the nucleus, and those of H_{2}O, CO, CH_{4}, and C_{2}H_{6} extend to the ends of the east-west slit, ∼ 15′′ or 1000 km in either direction. A preliminary analysis of three lines of H_{2}O (2_{02}-3_{03}, 2_{11}-3_{12}, and 2_{12}-3_{13}) measured on Mar. 24.5 suggests a rotational temperature ∼ 70 K within a 1′′ × 1′′.4 box centered on the nucleus. The flux measured for the 2_{02}-3_{03} line was 4.8 × 10^{−18} watt/m^{2} within that aperture, corresponding to Q(H_{2}O) = 1.5 × 10^{29} molecules/s. Two branches of the ethane ν_{7} band were detected on Mar. 24.5, and are extended along the slit; they appear peaked on the nucleus and exhibit the spatial profile expected for a parent volatile. The flux for the ^{r}Q_{0}-branch (2983.4 cm^{−1}) was 2.4 × 10^{−17} watt/m^{2}, and that for the ^{r}Q_{1}-branch was 2.8 × 10^{−17} watt/m^{2}, in a 1′′ × 1′′.4 box centered on the nucleus. Assuming fluorescence excitation of the ^{r}Q_{0} upper levels from the K = 0 and K = 2 lower levels, the combined production rate for molecules in those levels is 2.2 × 10^{26} molecules/s. Fluorescence excitation of the upper levels of ^{r}Q_{1} leads to a combined production rate for molecules in the K = 1 and K = 3 lower levels of 2.6 × 10^{26} molecules/s. Their sum (4.8 × 10^{26} molecules/s) is a lower bound to the total production rate for C_{2}H_{6} in K = 0-3. If the K-ladders exhibit a 130-K distribution, then the total production rate could be larger by ∼ 3. The mixing ratio for C_{2}H_{6} is 0.3 percent in the observed states alone.”

Naked-eye m_{1} estimates: Mar. 30.14 UT, 1.5 (M. V. Zanotta, Monte Generoso, Italy); 30.88, 1.7 (A. Diepvens, Balen, Belgium); 31.22, 1.5 (R. Keen, Mt. Thorodin, CO); 31.79, 1.9 (J. Tichá, Kleť, Czech Republic).

1996 April 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green